### v1

# revtex4-2Repair the float

Abstract
We report temperature measurements using a transmon qubit by detecting the population of the first three levels of it, after employing a sequence of $\pi$-pulses and performing projective dispersive readout. We measure the effective temperature of the qubit and characterize its relaxation and coherence times $\tau_{1,2}$ for three devices in the temperature range 203002030020–30020 – 300 mK. Signal-to-noise (SNR) ratio of the temperature measurement depends strongly on $\tau_{1}$, which drops at higher temperatures due to quasiparticle excitations, adversely affecting the measurements and setting an upper bound of the dynamic temperature range of the thermometer. The measurement relies on coherent dynamics of the qubit during the $\pi$-pulses. The effective qubit temperature follows closely that of the cryostat in the range $100-250$ mK. We present a numerical model of the qubit population distribution and compare it favorably with the experimental results.
preprint: APS/123-QED
I Introduction
In recent years theoretical proposals for quantum thermometry have attracted a lot of attention due to prospects of monitoring temperature with minimal invasion, properties of quantum systems with high accuracy at extremely low temperatures and opportunities for studying open quantum systems and quantum thermodynamics [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. However, the number of conceptually different experimental realizations of quantum thermometers based on coherent quantum objects remains quite limited. For example, quantum thermometry was demonstrated using a nuclear magnetic resonance (NMR) techniques for polarization measurements of spin ensembles in solutions at room temperature [17] and using coherent quantum N00N-states [18]. Generation and transfer of the latter was recently realized on the basis of a superconducting qubit [19]. Also there are noteworthy experiments on thermometry based on nitrogen-vacancy (NV) colour centres in diamond [20, 21, 22, 23, 24]. Another possible platform for quantum thermometers is semiconductor quantum dots, which were used for spectroscopy-type measurements in sub-kelvin temperature range in Ref. [25], exhibiting good signal-to-noise ratio and fast operation but limited coherence.
Prospects of utilizing superconducting qubits as thermometers are of dual research interest: they can facilitate studies of open quantum systems and interaction between the qubits and their environment, which is vital in the context of building large multiqubit systems such as quantum computers and simulators. On the other hand qubit thermometers can be used for realization of quantum thermometry protocols. Quantum electrodynamics (QED) circuits based on coherent superconducting transmon qubits were primarily used to measure the residual qubit population and the corresponding residual temperature in a number of works, for example in [26, 27, 28, 29]. The residual population of a transmon in a 3D cavity was measured by Rabi oscillations on the $e-f$ transition in Ref. [27] and correlation measurements in Ref. [28]. In Ref. [29] another technique of applying sequences of $\pi$-pulses was used to measure the effective qubit temperature by observing the population distribution for the three lowest energy states of a transmon in a planar geometry. A different approach was demonstrated in Ref. [30], where thermometry was done by reflection-type spectroscopy measurements of a transmon qubit strongly coupled to a waveguide.
Yet qubits are sensitive to various sources of dissipation and decoherence, and the diverse nature of their environment (see, for example, [31, 32, 33, 34, 35, 36, 37, 38, 39]). Therefore in practice, identification of a specific thermal bath, leading to thermalization of the qubit, can be quite challenging; i.e., it may be difficult to identify what temperature the qubit is actually measuring.
In this paper we present a comprehensive and concrete study of employing a superconducting qubit for quantum thermometry at sub-kelvin temperatures. The measurement technique used in our work relies on applying sequences of $\pi$-pulses for manipulating the qubit population distribution, requiring sufficient coherence of the qubit, and then performing consecutive qubit state readout, which was described in Ref. [29]. We also study temperature dependencies of qubit relaxation and coherence times. They determine the working range of the thermometer in terms of temperature, and they give conatraints on the pulse durations in the measurement. While showing a solid thermalization of the qubit to the cryostat stage, the technique allows us to measure the population distribution at temperatures up to $200-220$ mK, when the second excited state of the qubit was already significantly populated. This exceeds the applicability range of methods based on consideration of only the lowest qubit levels. At the same time the technique is less dependent on the signal-to-noise ratio (SNR) of the measurement setup. We compare the experimental results with a numerical model of qubit population measurements, being in a good qualitative agreement. Within the model we investigate the influence of the measured parameters on the population and effective qubit temperature obtained as the result. Finally, we address the factors limiting the dynamical range of a superconducting qubit thermometer. The paper is organized as follows: in Section II we provide a brief description of a realistic model of a qubit being in thermal equilibrium with heat baths. Section III explains the principle of the population distribution measurements. Section IV provides a brief description of the samples used in the experiment, and discussion on the main experimental results is given in Section V, and finally, limits of the thermometer and ways of mitigation are discussed in Section VI. The paper is followed by appendices providing additional experimental data, details of the numerical simulations, experimental setup and sample fabrication.
II Qubit population and effective temperature
Let us consider a qubit in thermal equilibrium with a heat bath having temperature $T$. In this case the qubit is in a thermal state, which means that its density matrix is diagonal and the probabilities $p_{g}$ and $p_{e}$, respectively, to be in the ground |$\ket{g}$ or excited state |$\ket{e}$ (qubit population) are Boltzmann distributed as
$p_{e}/p_{g}=e^{-\frac{\hbar\omega_{ge}}{k_{B}T}},$ (1)
where $\hbar\omega_{ge}$ is the separation between these two states, and $k_{B}$ is the Boltzmann constant. Thus, by measuring the population ratio of the qubit one obtains the temperature of the thermal bath.
In a more realistic scenario when a qubit is in steady state with respect to several uncorrelated heat baths (with temperatures $T^{(i)}$, where $i$ enumerates baths), the state of the qubit is still a thermal-like state. But the temperature characterizing this state is not necessarily the temperature $T^{(i)}$ of any of the thermal baths. Then Eq. (1) could be interpreted as giving an effective qubit temperature $T_{\mathrm{eff}}$. In this sense, Eq. (1) can be rewritten as a definition of $T_{\mathrm{eff}}$ as
$T_{\mathrm{eff}}=\frac{\hbar\omega_{ge}}{k_{B}}\left(-\ln{\frac{p_{e}}{p_{g}}}\right)^{-1},$ (2)
which coincides with the true temperature in equilibrium.
For the case of ohmic heat baths, one can find the energy relaxation rate $\gamma_{1}$ for the qubit towards $T_{\mathrm{eff}}$ using Fermi golden rule approach. The detailed description is given in Appendix A. The qubit energy relaxation rate writes
$\gamma_{1}=\sum_{i}\left(\Gamma_{\downarrow}^{(i)}+\Gamma_{\uparrow}^{(i)}\right)=\sum_{i}\gamma_{1}^{(i)}\left[2n\left(T^{(i)},\omega_{ge}\right)+1\right]\\ =\sum_{i}\gamma_{1}^{(i)}\coth{\frac{\hbar\omega_{ge}}{2k_{B}T^{(i)}}},$ (3)
where $n=\left[\exp{\left(\hbar\omega_{ge}/k_{B}T\right)}-1\right]^{-1}$ is the Bose-Einstein distribution.
The population ratio for the qubit then writes
$\frac{p_{e}}{p_{g}}=\frac{\Sigma_{i}\,\Gamma_{\uparrow}^{(i)}}{\Sigma_{i}\,\Gamma_{\downarrow}^{(i)}}=\frac{\Sigma_{i}\,\gamma_{1}^{(i)}n\left(T^{(i)},\omega_{ge}\right)}{\Sigma_{i}\,\gamma_{1}^{(i)}\left[n\left(T^{(i)},\omega_{ge}\right)+1\right]}$ (4)
III Population measurements
In this work we measured population distribution of several transmon qubits to obtain temperature using the technique described in [29], which is based on applying different sequences of $\pi$-pulses and then performing a dispersive readout of the qubit state. Here we do not assume that the transmon is a two-level qubit, but include the higher levels, especially the third one |$\ket{f}$ in the analysis and protocols. We presume, that the population of a qubit thermalized at temperature $T$ follows the Maxwell-Boltzmann distribution $p_{i}=\exp{\left(-E_{i}/k_{B}T\right)}/Z$, with the partition function $Z=\sum_{i=1}^{3}\exp{\left(-E_{i}/k_{B}T\right)}$ and the $i$-th level energy $E_{i}$. A transmon qubit typically has a weak anharmonicity $\alpha$ $(|\alpha|\ll\hbar\omega_{ge})$ [40], so that the energy separations $\hbar\omega_{ge}$ and $\hbar\omega_{ef}$ are related by $\omega_{ef}=\omega_{ge}+\alpha/\hbar\cong\omega_{ge}$. For example, if a transmon with $\omega_{ge}/2\pi=6.65\,$GHz and $\alpha/h=-230\,$MHz is thermalized at temperature 300300300\,300mK we can estimate the population ratios $p_{e}/p_{g}\approx 0.35$ and $p_{f}/p_{g}\approx 0.12$. Thus, it is important to take into account the finite population of the $f$-level. Limiting to the three lowest energy states of a transmon, the density matrix of a thermalized transmon reads
[…]
Let us denote the readout response of a projective measurement of the qubit being in state |$\ket{i}$ as $\varphi_{i}$, which we will call “pure state response” further in the text. Then the result $\varphi$ of the qubit state measurement is a linear combination of the form $\varphi=p_{g}\varphi_{g}+p_{e}\varphi_{e}+p_{f}\varphi_{f}$. The $\pi_{ge}$($\pi_{ef}$)-pulses effectively swap pairs of the population probabilities $p_{g}$ and $p_{e}$ ($p_{e}$ and $p_{f}$), which results in getting different readout outcomes, which we denote $x_{i}$ and $y_{j}$, $i,j=0,1,2$. An example of such a measurement is shown in Fig. 1 a, c. We can then write a system of linear equations as
$x_{0}=p_{g}\varphi_{g}+p_{e}\varphi_{e}+p_{f}\varphi_{f},\quad\quad\left(\mathrm{no\ pulse}\right)$
$x_{1}=p_{e}\varphi_{g}+p_{g}\varphi_{e}+p_{f}\varphi_{f},\quad\quad\quad\left(\pi_{ge}\right)$
$x_{2}=p_{e}\varphi_{g}+p_{f}\varphi_{e}+p_{g}\varphi_{f},\quad\quad\ \left(\pi_{ge}\pi_{ef}\right)$ (6)
$y_{0}=p_{g}\varphi_{g}+p_{f}\varphi_{e}+p_{e}\varphi_{f},\quad\quad\quad\left(\pi_{ef}\right)$
$y_{1}=p_{f}\varphi_{g}+p_{g}\varphi_{e}+p_{e}\varphi_{f},\quad\quad\ \left(\pi_{ef}\pi_{ge}\right)$
$y_{2}=p_{f}\varphi_{g}+p_{e}\varphi_{e}+p_{g}\varphi_{f}.\quad\quad\left(\pi_{ef}\pi_{ge}\pi_{ef}\right)$
Figure 1: Scheme of the experiment and the sample design. a) The pulse sequence used in the $x_{2}$-measurement. b) A scheme of the qubit operations, depicting the numerical simulation of the $x_{2}$-measurement. c) A schematic depiction of how the $x_{2}$-sequence of $\pi$-pulses changes the qubit population distribution. Initially the qubit population corresponds to the thermalized state. Then the $\pi$-pulses consecutively swap populations of the pairs of qubit levels. Finally, there is a finite length readout duration, during which the population is still evolving. d) An optical image of the transmon and readout resonator coupled to the waveguide (sample Q2-III). e) An optical image of the transmon, shown on panel d). f) SEM image of the Al/AlOx/Al tunnel contact of the transmon, circled in panel e).
We can find functional dependencies between $p_{i},i\in\{g,e,f\}$ and $x_{j},y_{k}$ with $j,k=0,1,2$ which are valid for arbitrary pure state responses $\varphi_{i}$. These quantities are temperature dependent given by $A=\frac{p_{g}-p_{e}}{p_{g}-p_{f}}$, $B=\frac{p_{e}-p_{f}}{p_{g}-p_{e}}$ and $C=\frac{p_{e}-p_{f}}{p_{g}-p_{f}}$, which can also be expressed through the readout results $x_{j},\ y_{k}$ (see Table 2). To understand the temperature behaviour of functions $A,\ B$ and $C$, we can take the low temperature limit $\hbar\omega_{ge}\ll k_{B}T$, also assuming $p_{f}\rightarrow 0$. Then we can rewrite $A\approx 1-p_{e}/p_{g}$, $B\approx C\approx p_{e}/p_{g}$ or
$A\approx 1-$ $\exp{\left(-\hbar\omega_{ge}/k_{B}T\right)},$ (7)
$B\approx C\approx$ $\exp{\left(-\hbar\omega_{ge}/k_{B}T\right)}.$ (8)
IV Sample description
We used transmon qubits on three different chips I, II and III (which is reflected in the name, e.g. R2-I, R4-I, R3-II, Q2-III) made in different fabrication cycles. All transmon qubits had a fixed transition frequency $\omega_{ge}/2\pi\approx 6.5\,$GHz or flux-tunable spectrum ($\omega_{ge}/2\pi=5.5-6.8\,$GHz) and charging energy $E_{c}/h=210-230\,$MHz (see Fig. 1) and they were weakly coupled ($g/2\pi=35-45$ MHz) to a coplanar waveguide (CPWG) $\lambda/4$ readout resonator with the internal quality factor $Q_{\mathrm{int}}\sim 10^{5}$, external quality factor $Q_{\mathrm{ext}}\approx 8\cdot 10^{4}$ and fundamental frequency $\omega_{r}/2\pi=4.5-6.0\,$GHz for a dispersive state readout ($\left|\Delta_{r}\right|=\left|\omega_{ge}-\omega_{r}\right|\gg g$). Experimentally found parameters of the devices are shown in Table 1. The fabrication details are given in Appendix D.
Table 1: Parameters of the devices.
Device $\omega_{ge}/2\pi$, GHz $\omega_{ef}/2\pi$, GHz $E_{c}/h$, MHz $\Delta_{r}/2\pi$, GHz $g/2\pi$, MHz
R2-I 6.422 6.221 201 1.876 34
R4-I 6.649 6.417 232 1.753 38
R3-II 6.732 6.513 219 0.765 44
Q2-III 7.042 6.835 207 2.151 37
V Results and discussion
In the experiment we measured qubit relaxation and coherence times $\tau_{1}$ and $\tau_{2}$, and qubit effective temperature $T_{\mathrm{eff}}$, while controlling the temperature of the mixing chamber (MXC) stage of the dilution refrigerator. After a certain MXC stage temperature $T_{\mathrm{MXC}}$ was set, we waited for 10 min for thermalization of the sample stage and then performed a series of measurements, consisting of five repetitions of the following measurement sequence: (i) measurement of the MXC stage temperature $T_{\mathrm{MXC}}$, (ii) population distribution measurement $x_{i},\ y_{j}$, (iii) measurement of qubit relaxation time $\tau_{1}$, (iv) Ramsey oscillations measurement to get $\tau_{2}^{R}$ and $\omega_{ge}$ and (v) Hahn echo measurement with the decay time $\tau_{2}^{E}$. Each of these iterations took approximately 5 min. Qubit excitation pulses were applied via an AC drive antenna, and the readout pulses were sent through a waveguide. The measurement setup is shown in Fig. 8.
First we address the temperature dependence of the qubit decay time $\tau_{1}$, shown in Fig. 2 a. It stays almost constant in temperature range $20-150$ mK. Higher than $T_{\mathrm{MXC}}\approx 170\,$mK a noticeable effect of quasiparticle induced relaxation appears and already at $T_{\mathrm{MXC}}\approx 250-270\,$mK decay time drops to below $0.5\,\mu$s, which is shorter than the readout pulse duration $\Delta t_{\mathrm{RO}}=1.5-2.0\,\mu$s and becomes comparable to the duration of $\pi$-pulse sequences used for the population measurements. To compare $\tau_{1}$ with theoretical prediction we use the quasiparticle relaxation model in Ref. [41] that gives the relaxation rate for a transmon as
$\begin{split}\gamma_{1}^{qp}=\frac{1}{\pi}&\frac{\omega_{p}^{2}}{\omega_{ge}}\left\{x_{qp}\sqrt{\frac{2\Delta}{\hbar\omega_{ge}}}\right.\\ &+\left.4e^{-\frac{\Delta}{k_{B}T}}\cosh\left(\frac{\hbar\omega_{ge}}{2k_{B}T}\right)\,K_{0}\left(\frac{\hbar\omega_{ge}}{2k_{B}T}\right)\right\}\\ \end{split}$ (9)
with the plasma frequency $\omega_{p}=\sqrt{8E_{J}E_{C}}/\hbar$, equilibrium quasiparticle density $x_{qp}=\sqrt{2\pi k_{B}T/\Delta}\,\exp\left(-\Delta/k_{B}T\right)$, modified Bessel function of the second kind $K_{0}$ and the superconducting gap $\Delta$. Expression (9) provides a good correspondence with the experimentally observed drop of $\tau_{1}$ at temperatures above 220 mK for the gap value $\Delta_{\mathrm{Al}}/e=180\,\mu\mathrm{V}$ (see Fig. 2 b). Thus, the sharp exponential drop of quasiparticle-induced suppression of $\tau_{1}$ restricts the measurement range for Al-based qubits using long pulse sequences. This is clearly demonstrated in the experimental data discussed below.
[…]
As to the dephasing during the $\pi$-pulses, typical length $\Delta t_{\pi}$ of single $\pi$-pulses varied between 20 and 220 ns, with the lower limit $\Delta t_{\mathrm{min}}$ determined either by the qubit anharmonicity ($\Delta t_{\mathrm{min}}>h/|\alpha|\simeq 1/200\,$MHz $=5\,$ns) or by the dynamical range of instruments and total line attenuation of the setup (see Fig. 8). We present experimental data of the temperature dependence of the dephasing time $\tau_{\varphi}$ in Fig. 2 b. The times $\tau_{\varphi}^{R}$ and $\tau_{\varphi}^{E}$ are extracted from the measurements of Ramsey oscillations and Hahn echo measurements, respectively, as $\tau_{\varphi}^{R,E}=1/\gamma_{\varphi}^{R,E}=1/\left(\gamma_{2}^{R,E}-\gamma_{1}/2\right)$, where the decoherence rates $\gamma_{2}^{R,E}$ describe Ramsey oscillations and Hahn echo experiments. While in the low-temperature limit $\tau_{\varphi}^{R,E}\gtrapprox\tau_{1}$, with $\tau_{\varphi}^{R}$ exhibiting a clearly visible saturation below $50-70$ mK, the dephasing time starts decreasing at lower temperatures than $\tau_{1}$ and reaches $1-2\,\mu$s at 150 mK, becoming difficult to measure at $220-250\,$mK due to the finite ring-up time of the readout resonator $\sim Q_{ext}/\omega_{r}\simeq 1\,\mu$s. Similarly to the analysis of the relaxation time, we can estimate influence of quasiparticles on the dephasing, using the approach presented in Ref. [33]. The pure dephasing rate due to tunneling of equilibrium quasiparticles is given by expression
$\gamma_{\varphi}=\frac{E_{c}}{\pi\hbar}\frac{k_{B}T}{\Delta}\exp{\left(-\frac{\Delta}{k_{B}T}\right)},$ (10)
and appears to be negligible in the experimental temperature range. However, we can estimate the dephasing mechanism caused by the Josephson frequency fluctuations due to quasiparticles occupying the Andreev bound states in the tunnel junction of the transmon. The latter writes [33]
$\gamma_{\varphi}\sim 4\pi\frac{\omega_{p}^{2}}{\omega_{ge}}\sqrt{\frac{x_{qp}^{A}}{N_{e}}},$ (11)
[…]
Population ratios $A,\ B$ and $C$ can be extracted from the measurement data $x_{i},\ y_{j}$ by three different methods each, which we denote as $A_{1},A_{2},A_{3},B_{1},\dots,C_{3}$ (see Tab. 2). Qualitatively speaking, the difference between the methods is the order of the $\pi$-pulses in the control sequence. Because of the qubit relaxation and relatively low population of the $f$-level, the resulting signals can have different signal-to-noise (SNR) ratio, giving us some guidance on how to choose the best one (see Fig. 5 and discussion in Sec. VI and Appendix C).
An example of population measurements in terms of datasets $A_{1},A_{2},\dots C_{3}$ (see exact expressions in Appendix in Table 2) for device R4-I can be seen in Fig. 2 c. While according to expressions (7), (8), in the low-temperature limit $A\rightarrow 1$ and $B,C\rightarrow 0$, there is a clear saturation at $p_{e}/p_{g}\approx 0.024$, which agrees with $T_{\mathrm{eff}}\approx 85\,$mK, calculated by Eq. (2). Above 170 mK the population readings have increased deviation from the analytical expressions caused by the enhanced relaxation.
Measurements of the qubit effective temperature for devices R2-I, R4-I and Q2-III extracted from $A_{2}$ are shown in Fig. 2 d and demonstrate a linear behaviour in the temperature range $120-230$ mK close to one-to-one correspondence with the temperature of the mixing chamber stage $T_{\mathrm{MXC}}$ of the cryostat. At lower temperatures of the cryostat $T_{\mathrm{eff}}$ exhibits saturation, reaching $T_{\mathrm{eff}}^{0}\approx 65-85\,$mK, which coincides with the saturation temperature estimated from the population in Fig. 2 c. This elevated temperature compared to $T_{\mathrm{MXC}}$ is most probably caused by imperfect shielding against thermal photons and insufficient filtering of the measurement lines. For comparison, $T_{\mathrm{eff}}$ extracted using all nine methods $A_{1},A_{2},A_{3},B_{1},\dots,C_{3}$ are shown in Fig. 3 a. A thorough explanation of the deviations between the experimental data and the expected behaviour above 170 mK in Figs. 2 c, d requires consideration of the time evolution of the qubit population due to its finite lifetime, which we discuss in the following subsection.
V.1 Comparison of the Experimental Data with the Numerical Model of the Qubit Population Evolution
In order to understand the deviation of the population functions $A,B$ and $C$ from the exact formulas in Tab. 2 and the qubit effective temperature $T_{\mathrm{eff}}$ from the cryostat temperature $T_{\mathrm{MXC}}$ at higher temperatures, we implement a numerical model which accounts for time evolution of the population distribution of a qubit. First, let us discuss important assumptions underlying the model, which are based on the experimental data. We assume, that the control sequences of $\pi$-pulses are short compared to the qubit dephasing and relaxation times $\left(3\Delta t_{\pi}<\tau_{\varphi},\tau_{1}\right)$. Then we are interested in tracking the time evolution of the population distribution, described by the diagonal terms of the density matrix. The duration of the qubit state readout $\Delta t_{\mathrm{RO}}$ is larger than or of the order of the relaxation time $\left(\Delta t_{\mathrm{RO}}\gtrsim\tau_{1}\right)$. The model considers population of the three lowest states with finite decay times $\tau_{1}^{ge}$ and $\tau_{1}^{ef}$ of the $ge$- and $ef$-transitions, respectively.
We begin with a qubit, which is thermalized at a temperature $T$, setting initial populations $p_{g,e,f}^{(0)}(T)$ and then imitate application of $\pi_{ge}$- and $\pi_{ef}$-pulses by instantaneous swapping of the corresponding level populations with finite efficiencies, designated as $\delta_{ge,ef}\in[0,1]$ (see Fig. 1 e). For example, a $\pi_{ge}$-pulse with $\delta_{ge}=0.9$ means that a qubit state with $p_{g}=1.0,\ p_{e}=0.0,\ p_{f}=0.0$ would turn result into $p_{g}=0.1,\ p_{e}=0.9,\ p_{f}=0.0$. Then we calculate the time evolution of the population during the readout time, and multiply it by arbitrary pure state responses $\varphi_{i}$. Next we construct the population ratios $A,\ B$ and $C$ and calculate the effective temperature (see Fig. 3 b). A detailed description of the model is provided in Appendix B.
The numerical model can qualitatively reproduce the deviation of the qubit population functions $A,B$ and $C$ from the analytical expressions (see Fig. 6 in App. B), as well as the resulting deviations of the temperatures $T_{\mathrm{eff}}$ extracted with different methods in the higher temperature range, which is presented in Fig. 3. This divergence is a consequence of the temperature-dependent qubit decay times $\tau_{1}^{ge}$ and $\tau_{1}^{ef}$, caused by quasiparticles. For the simulation results presented in Fig. 3 b, we use frequencies $\omega_{ge}$ and $\omega_{ef}$ and the temperature dependent relaxation time $\tau_{1}^{ge}(T)=2\pi/\left(\gamma_{1}^{qp}(T)+\gamma_{1}^{0}\right)$, which correspond to the sample R4-I shown as the red line in Fig. 2 b, and $\tau_{1}^{ef}\approx\tau_{1}^{ge}/2$, valid for a transmon qubit [40]. For ideal $\pi$-pulses, the effective temperatures $T_{A1},T_{A2},\dots,T_{C3}$ extracted from different ratios $A_{1},A_{2},\dots,C_{3}$, respectively, demonstrate almost one-to-one behaviour for the temperature range $30-200$ mK, when the qubit lifetimes are relatively long. But above 200 mK, when $\tau_{1}$ is significantly suppressed, there is an increasing deviation between these dependencies. Importantly, this does not indicate lack of thermalization of the qubit, rather than showing the limits of applicability of the given thermometry protocol. Moreover, in this temperature limit the values $T_{A1},T_{A2},\dots,T_{C3}$ start to depend on the pure state responses $\varphi_{i}$, which eliminates the main advantage of the protocol. In fact, Eqs. (III) are no longer valid, as the qubit population changes significantly during the readout pulse.
Another error in determining the effective temperature of the qubit is the finite efficiency of the pulses. As seen in Fig. 3 b, the divergence of $T_{A1},T_{A2},\dots,T_{C3}$ grows monotonously with the temperature and is noticeable already at $100-150$ mK, which is a characteristic of the protocol. The plot, showing the relative error between the extracted temperature and the qubit thermalization temperature averaged over different methods $A_{1},A_{2},\dots,C_{3}$ is shown in Fig. 7 in Appendix B.
Figure 3: Effective qubit temperature extracted from the population distribution measurements. a) Experimental data for the qubit R4-I for all nine ratios $A_{1},A_{2},A_{3},B_{1},\dots,C_{3}$. b) Effective qubit temperature, extracted from the qubit population distribution according to the numerical model (see details in main text). Using the device parameters of R4-I listed in Table 1, we initialize the qubit population $p_{i}(T),i\in\{g,e,f\}$ Boltzmann-distributed at temperature $T_{\mathrm{qubit}}$ and imitate the measurement procedure. We consider the temperature dependence of qubit lifetime $\tau_{1}$, which we approximate from experimental data shown in Fig. 2. Solid lines show the result of simulating perfect $\pi$ pulses, which we compare to finite pulse efficiency $\delta_{ge}=0.9,\delta_{ef}=0.8$ (dotted lines). The simulation was done with parameters $2\tau_{1}^{ef}(T_{\mathrm{qubit}}\rightarrow 0)=\tau_{1}^{ge}(T_{\mathrm{qubit}}\rightarrow 0)=5.5\,\mu s$, $\pi$-pulse duration is 165 ns and readout pulse length $\Delta t_{\mathrm{RO}}=2\,\mu$s.
V.2 Qubit Thermalization
In order to make sure, that the qubits used in our measurements thermalize to a single thermal bath, which is the substrate, we analyze the relaxation rate $\gamma_{1}$ and photon number $n\left(\omega_{ge},T\right)$ as functions of the qubit effective temperature $T_{\mathrm{eff}}$. For example, if a qubit is in a steady state with several ohmic baths, then from (3) and (25) we can write down the following temperature dependencies for $\gamma_{1}$ and $n\left(\omega_{ge},T_{\mathrm{eff}}\right)$:
$\gamma_{1}(T$ $)=\gamma_{1}^{0}\left[2\,n(\omega_{ge},T_{\mathrm{eff}})+1\right]=\gamma_{1}^{0}\,\coth{\frac{\hbar\omega_{ge}}{2k_{B}T_{\mathrm{eff}}}},$ (12)
$n(\omega_{ge},T_{\mathrm{eff}})=\frac{\gamma_{1}^{0}}{\sum_{i}\gamma_{1}^{(i)}}n\left(\omega_{ge},T_{\mathrm{MXC}}\right)+n_{0},$ (13)
where we denote the relaxation rate and the residual qubit population at the base temperature as $\gamma_{1}^{0}$ and $n_{0}$, respectively. The corresponding experimental data is presented in Fig. 4. The photon number dependence of $\gamma_{1}$ (see Fig. 4 a) exhibits two regimes: the linear growth described by Eq. (12), where the derivative $\partial\gamma_{1}/\partial n$ is twice bigger than the offset $n_{0}$, and the sharp rise due to the quasiparticle relaxation.
The dependence between the photon number $n\left(\omega_{ge},T_{\mathrm{eff}}\right)$, which corresponds to the qubit effective temperature, and $n\left(\omega_{ge},T_{\mathrm{MXC}}\right)$, defined by the cryostat temperature, characterizes the general process of the qubit thermalization (see Fig. 4 b). In case of a single thermal bath coupled to the qubit, the dependence would be exactly one-to-one, whereas the presented data shows finite photon population $n_{0}$ in the low-temperature limit and the slope slightly less than unity, witnessing coupling to other thermal baths. One of those could be a residual radiation field, as well as a bath of TLSes or nonequilibrium quasiparticles with the effective temperature $65-85$ mK, which sets the lower limit for the population of the qubit at the lower temperatures of the cryostat.
Figure 4: a) Measurements of the qubit relaxation rate $\gamma_{1}$ as a function of the photon number $n\left(\omega_{ge},T_{\mathrm{eff}}\right)$ for qubits R2-I and Q2-III. The solid lines are linear fits $y=kx+b$, with parameters $k=0.71\pm 0.23$ MHz/photon and $b=0.41\pm 0.07$ MHz for R2-I and $k=0.63\pm 0.39$ MHz/photon and $b=0.37\pm 0.09$ MHz for Q2-III, respectively, in line with Eq. (12), i.e. $k\simeq 2b$. The dashed lines show the quasiparticle-induced relaxation $\gamma_{1}^{qp}$ according to Eq. (9). b) Dependence between the photon numbers $n\left(\omega_{ge},T_{\mathrm{eff}}\right)$ and $n\left(\omega_{ge},T_{\mathrm{MXC}}\right)$. The dashed line shows one-to-one correspondence and the solid lines are linear fits with parameters $k=0.965\pm 0.114$ and $b=0.017\pm 0.037$ for R2-I and $k=0.900\pm 0.140$ and $b=0.012\pm 0.045$ for Q2-III, respectively. Here, $k=\gamma_{1}^{0}/\sum_{i}\gamma_{1}^{(i)}\approx 1$ (see Eq. 13) means that thermalization of the qubit is dominated by its coupling to the thermal bath with the temperature $T_{\mathrm{MXC}}$.
VI Thermometer Limitations and Ways of Improvement
VI.1 Temperature Range
The Al-based transmon thermometers studied in this work showed a relatively narrow working temperature range $60-220$ mK. In this subsection we discuss possibilities of widening this regime.
A higher residual temperature of the qubits $T_{\mathrm{eff}}=65-85$ mK compared to the base cryostat temperature $T_{\mathrm{base}}\approx 20$ mK is most likely caused by both improper shielding of the sample stage and filtering of the measurement lines [43], likely due to nonequlibrium quasiparticles [27] or TLSes [28, 34]. Recent works demonstrated the possibility of achieving effective qubit temperatures of $30-45$ mK [27, 28, 30, 44].
When considering the higher temperature limit of the presented method the following aspects should be taken into account: i) consideration of only the three lowest states is valid until the population of the next $d$-level is negligible, otherwise, at higher temperatures, the system (III) should be supplemented with appropriate terms having $\varphi_{d}$ and $p_{d}$; ii) suppression of the qubit relaxation time $\tau_{1}$ due to growing quasiparticle density above 200 mK; iii) sufficiently long coherence time for the qubit state preparation. In agreement with the model, quasiparticle suppression of $\tau_{1}$ leads to a significant influence on the population measurements, which in turn affects adversely on determining the effective temperature. This is a hard limit due to the exponential growth of the quasiparticle density with temperature.
Another possibility is employment of materials with a larger superconducting gap to shift the quasiparticle suppression limit to higher temperatures. This might be done by using Nb-based tunnel junctions [45, 46, 47, 48] with the critical temperature $T_{c}^{\mathrm{Nb}}=7.6-8.7$ K or by replacing the tunnel junction with a nanowire made of granular Al having critical temperatures $T_{c}^{\mathrm{grAl}}$ up to 3.15 K [49, 50, 51], which is 2.5 times larger than critical temperature of bulk aluminium, $T_{c}^{\mathrm{Al,bulk}}=1.2$ K.
The qubit state preparation is also of big importance in the population distribution measurement presented here. In this sense, the qubit coherence time, or more strictly speaking, the dephasing time should be longer than the control pulse sequences $\left(\tau_{\varphi}\gtrsim 3\Delta t_{\pi}\right)$. Since the dephasing rate is growing with the temperature increase, it limits duration of the pulse sequences at higher temperatures.
VI.2 Signal-to-Noise Ratio and Accuracy
The parameters of the experimental setup can be fine-tuned to improve both the temperature measurement accuracy and the range of the cryostat temperatures, so that it reaches the quasiparticle limit. First of all, improvement of the $\pi$-pulse efficiency affects positively on the measurement accuracy. Secondly, as mentioned above, the short lifetime of the qubit is the main restriction of the method. We can divide the qubit state decay into two parts: i) decay during the control sequence and ii) decay during the readout. The first process leads to change of the qubit population distribution $p_{i}$ and can be associated to imperfect control pulses, leading to increased deviation between $T_{A1},T_{A2},\dots,T_{C3}$. Decay during the readout pulse is defined mainly by the ring-up time of the resonator, and diminishes distinguishability between the qubit states (i.e. difference between $\varphi_{i}$ on the $I-Q$ plane), thus reducing the SNR, and also distorting the extracted effective temperature. The main method of improving the SNR is optimizing the ratio $\chi/\kappa$ for the readout resonator, where $\chi$ is the dispersive shift and $\kappa$ is the linewidth, as well as an appropriate readout frequency detuning for a given $\chi/\kappa$, which directly affects the pure state responses $\varphi_{i}$ [52].
Another factor directly affecting the efficiency of the $\pi$-pulses is temperature dependent properties (specifically S21) of the measurement lines of the cryostat and CPWGs on the chip. In particular, it is known, that superconducting CPWG resonators have a temperature dependent Q-factors and fundamental frequencies [53, 54, 55, 56, 57, 58, 59], which can lead to changes in amplitude of the qubit drive signal. Moreover, the qubit frequency even for the single-junction transmons can slightly vary with temperature (less than 50 kHz in our case). All these factors can lead to an increased deviation of the $\pi$-pulse parameters if a calibration procedure is not applied at every cryostat temperature point.
The minimal temperature measurement error is defined by the Cramer-Rao bound from the quantum Fisher information (QFI) [5], where we can use an expression for the relative error $\left|\Delta T/\langle T\rangle\right|_{\mathrm{sm}}$ of a single measurement for a two-level thermometer with energy separation $\hbar\omega_{ge}$ and $N$-fold degenerate excited level as
$\left(\frac{\Delta T}{\langle T\rangle}\right)^{2}_{\mathrm{sm}}=\frac{\left(N-1+e^{x_{ge}}\right)^{2}}{(N-1)x_{ge}^{2}e^{x_{ge}}},$ (14)
$x_{ge}=\hbar\omega_{ge}/k_{B}T$. An analogous expression can be derived for the three-level case considered in our work,
$\left(\frac{\Delta T}{\langle T\rangle}\right)^{2}_{\mathrm{sm}}=\frac{\left(e^{x_{ge}+x_{gf}}+e^{x_{ge}}+e^{x_{gf}}\right)^{2}}{\left[x_{ge}^{2}e^{x_{gf}}+x_{gf}^{2}e^{x_{ge}}+\left(x_{ge}+x_{gf}\right)^{2}\right]e^{x_{ge}+x_{gf}}},$ (15)
here $x_{gf}=\hbar\omega_{gf}/k_{B}T$ and $\omega_{gf}=2\omega_{ge}-|\alpha|$. Taking into account averaging over $N=2^{17}$ repetitions, one gets the lower bound for the relative error $\left(\Delta T/\langle T\rangle\right)^{2}=\left(\Delta T/\langle T\rangle\right)^{2}_{\mathrm{sm}}/N\approx 5.0\cdot 10^{-5}$ or $\left|\Delta T/\langle T\rangle\right|\approx 0.7\%$ for $\langle T\rangle=65$ mK and qubit frequency $\omega_{ge}=7.04$ GHz. The noise equivalent temperature $\mathrm{NET}=\sqrt{\left|\Delta T\right|^{2}\Delta t_{\mathrm{meas}}}$ with the measurement time $\Delta t_{\mathrm{meas}}=29$ s for our setup in the QFI-limited case would be $\mathrm{NET}_{\mathrm{QFI}}\approx 2.5$ mK$/\sqrt{\mathrm{Hz}}$. These estimates are an order of magnitude lower than the experimental errors. The data shown in Fig. 5 a has the relative errors $\left|\Delta T_{A}/\langle T_{A}\rangle\right|=14-42\%$, $\left|\Delta T_{B}/\langle T_{B}\rangle\right|=8.6-9.4\%$ and $\left|\Delta T_{C}/\langle T_{C}\rangle\right|=8.7-9.4\%$ at the base cryostat temperature $T_{\mathrm{MXC}}=22$ mK, which corresponds to $\mathrm{NET}\approx 28$ mK$/\sqrt{\mathrm{Hz}}$.
Figure 5: a) Time traces of the effective qubit temperature for the sample Q2-III at the base cryostat temperature $T_{\mathrm{MXC}}\approx 22\,$mK (shown as the black dashed line). b) Relative error of a single temperature measurement as a function of temperature. The symbols denote experimental time averaged data obtained from time traces of three samples: R2-I (green), R3-II (blue) and Q2-III (red). The error bars show the standard deviation. The curves denote the minimal error (i.e. the red filled area cannot be realized), defined by QFI for a two-level system (red dashed line) from the Eq. (14) for $N=2$, for a three level system with anharmonicity (solid dark red line) using Eq. (15), and for a $N$-level system with $N-1$ times degenerate excited level (red dotted line), calculated using Eq. (14) with $N=10$. The energy separations correspond to those defined for the sample Q2-III. The black dotted line shows the relative error for a three-level transmon with $\omega_{ge}/2\pi=3.00\,$GHz and $\omega_{ef}/2\pi=2.763$ GHz with the exponential rise shifted towards lower temperatures by $10-15$ mK.
Notably, if the thermometer has a constant level separation, the relative error grows exponentially at $T\ll\hbar\omega_{ge}/k_{B}$ (see Fig. 5), which is, however, not observed experimentally in our case, as the qubit effective temperature saturates at $T_{\mathrm{eff}}^{0}\approx 65-85$ mK. Our system is a three-level quantum thermometer with a constant anharmonic spectrum, which refers to sub-optimal quantum thermometers [5]. Knowing the dependence of QFI on the spectrum of the system, it is possible to optimize it for a better performance in the experimentally realistic temperature range by choosing an appropriate level separation and anharmonicity.
VI.3 Operation Speed
First of all, the presented method refers to the equilibrium type of quantum thermometry, meaning that the population measurement is performed after the qubit thermalization to the bath, which happens on time scales not shorter than the relaxation time $\tau_{1}$. In the presented experiment the time delay between the consecutive measurements, which is necessary for the qubit relaxation after the readout and thermalization of the qubit, was $\sim 10\,\tau_{1}$, which takes up most of the measurement time, and defines the speed of the response of the system. Pulse application and the qubit state readout can be limited to sub-microsecond range. Finally, the total measurement time is defined by the averaging number $t_{total}\sim N_{av}\cdot 10\,\tau_{1}=10^{6}\tau_{1}$, with $N_{av}\simeq 10^{5}$, meaning a time scale of seconds or minutes per point, depending on $\tau_{1}$.
VII Conclusion
In this work we experimentally realized equilibrium quantum thermometry of the cryostat temperature, based on the population distribution measurements of the three lowest energy levels of a transmon qubit. The presented technique allowed for temperature measurements in the range of $60-220\,$mK, in which the qubit population follows the Maxwell-Boltzmann distribution. Analysis of temperature dependencies of the qubit relaxation rate and the photon number showed that the qubit was thermalized to the cryostat. The working temperature range was defined by saturation of the effective qubit temperature in the low temperature limit and the quasiparticle suppression of the qubit relaxation time above $200$ mK. We compared the experimental data to a numerical model, describing time evolution of the qubit population, showing good agreement. We believe, a similar approach could be exploited for quantum thermometry of various on-chip structures and for studies of coupling between quantum objects to different thermal baths. Finally, this experimental platform allows for exploitation of quantum thermometry algorithms including nonequilibrium techniques for faster measurements and dynamical control of the thermometers for enhanced precision.
Acknowledgements.
The authors are grateful to Gershon Kurizki for fruitful discussions and Yu-Cheng Chang for assistance with measurements and fabrication. We acknowledge the provision of facilities and technical support by Aalto University at OtaNano - Micronova Nanofabrication Centre and OtaNano - Low Temperature Laboratory.
Appendix A Relaxation and Excitation Rates for Coupling to a Thermal Bath
Let us consider a quantum oscillator based on a resonator with frequency $\omega_{0}=1/\sqrt{LC}$ and characteristic impedance $Z_{0}=\sqrt{L/C}=50\,\Omega$, where $L$ and $C$ are the inductance and capacitance of the resonator, respectively, to which we connect an ohmic bath with resistance $R$ and temperature $T$. Then the excitation and relaxation rates $\Gamma_{\uparrow,\downarrow}$ caused by the bath can be written as [60]
$\Gamma_{\uparrow,\downarrow}=\frac{Z_{0}}{R}\frac{\mp\omega_{0}}{1-\exp(\pm\hbar\omega_{0}\beta)},$ (16)
where $\beta=1/k_{B}T$ is the inverse temperature and $k_{B}$ is the Boltzmann constant. For a more general case the resistance $R$ should be replaced by the real part of the bath impedance $Z$. Thus, we can introduce the quality factor $Q=\mathrm{Re}\,Z/Z_{0}$ and the resonator linewidth $\gamma=\omega_{0}/Q$, giving us
$\Gamma_{\uparrow,\downarrow}=\gamma\frac{\mp 1}{1-\exp(\pm\hbar\omega_{0}\beta)}.$ (17)
This can be further shortened for the sake of notation brevity to
$\Gamma_{\downarrow}=\gamma\left[n\left(\omega,T\right)+1\right],$ (18)
$\Gamma_{\uparrow}=\gamma\,n\,\left(\omega,T\right),$ (19)
where $n=1/\left[1-\exp\left(-\hbar\omega_{0}\beta\right)\right]$ is the Bose-Einstein distribution. The total decay rate in this case would be
$\Gamma_{1}=\Gamma_{\downarrow}+\Gamma_{\uparrow}=\gamma\left[2n\left(\omega,T\right)+1\right].$ (20)
Analogous expressions can be readily applied to the qubit-bath interaction (see for example [61, 60]).
A.1 Effective Temperature of a Two-Level System
Let us consider a qubit with the ground state |$\ket{g}$ and excited state |$\ket{e}$, thermalized with bath with temperature $T$. Then the population ratio of the qubit levels should follow the detailed balance principle $p_{e}/p_{g}=e^{-\hbar\omega\beta}=\Gamma_{\uparrow}/\Gamma_{\downarrow}$, and [62]
$p_{g}=\frac{\Gamma_{\downarrow}}{\Gamma_{\downarrow}+\Gamma_{\uparrow}},\quad p_{e}=\frac{\Gamma_{\uparrow}}{\Gamma_{\downarrow}+\Gamma_{\uparrow}}.$ (21)
Also, we can write it in terms of the qubit polarization:
$\langle\sigma_{z}\rangle=-p_{g}+p_{e}=-\tanh{(\hbar\omega\beta/2)}.$ (22)
Energy relaxation rate is just a sum of the two rates:
$\gamma_{1}=\frac{2\pi}{\tau_{1}}=\Gamma_{\downarrow}+\Gamma_{\uparrow}.$ (23)
A.2 Transmon Qubit Coupled to Several Baths
Here we will extend the results from the previous section to the case when a quantum object interacts with a number of uncorrelated baths with different temperatures $T^{(i)}$ and coupling rates $\Gamma_{\uparrow,\downarrow}^{(i)}$. As in previous section, the only requirement for the baths is that the corresponding rates obey the detailed balance principle. In this case we can just sum up the rates when calculating the level population:
$p_{g}=\sum_{i}\Gamma_{\downarrow}^{(i)}/\sum_{i}\left(\Gamma_{\downarrow}^{(i)}+\Gamma_{\uparrow}^{(i)}\right),$
$p_{e}=\sum_{i}\Gamma_{\uparrow}^{(i)}/\sum_{i}\left(\Gamma_{\downarrow}^{(i)}+\Gamma_{\uparrow}^{(i)}\right),$
$\frac{p_{e}}{p_{g}}=\frac{\Sigma_{i}\Gamma_{\uparrow}^{(i)}}{\Sigma_{i}\Gamma_{\downarrow}^{(i)}}=\frac{\Sigma_{i}\gamma_{1}^{(i)}n\left(T^{(i)},\omega\right)}{\Sigma_{i}\gamma_{1}^{(i)}\left[n\left(T^{(i)},\omega\right)+1\right]}$
The qubit energy relaxation rate writes
$\gamma_{1}=\sum_{i}\left(\Gamma_{\downarrow}^{(i)}+\Gamma_{\uparrow}^{(i)}\right)=\sum_{i}\gamma_{1}^{(i)}\left[2\,n\left(\omega,T^{(i)}\right)+1\right].$ (24)
One can use the same considerations for a resonator (harmonic oscillator). Then again the detailed balance principle should be applied to get $p_{n+1}/p_{n}=\sum_{i}\Gamma_{\uparrow}^{(i)}/\sum_{i}\Gamma_{\downarrow}^{(i)}$, which is positive, less then unity and does not depend on $n$. Thus we can find such a number $\beta_{\mathrm{eff}}=1/k_{B}T_{\mathrm{eff}}$, to have $p_{n+1}/p_{n}=\exp{\left(-\hbar\omega\beta_{\mathrm{eff}}\right)}$, leading to
$\bar{n}=\frac{\sum_{i}\Gamma_{\uparrow}^{(i)}}{\sum_{i}\left(\Gamma_{\downarrow}^{(i)}-\Gamma_{\uparrow}^{(i)}\right)}=\frac{\sum_{i}\gamma_{i}n\left(T^{(i)},\omega\right)}{\sum_{i}\gamma_{i}}=n\left(T_{\mathrm{eff}},\omega\right).$ (25)
Effective resonator temperature is then dependent on the number of photons $\bar{n}$:
[…]
All considerations above can be generalized to the case of an arbitrary number of baths.
Table 2: Different ways of the qubit effective temperature calculation from the qubit state readout results.
Ratio Calculation method Result Equation
1 2 3
A $\frac{x_{0}-x_{1}}{y_{0}-y_{1}}$ $\frac{y_{0}-x_{2}}{x_{0}-y_{2}}$ $\frac{y_{1}-y_{2}}{x_{1}-x_{2}}$ $\frac{p_{g}-p_{e}}{p_{g}-p_{f}}$ $A=\frac{1-\exp{\left(-\hbar\omega_{ge}/k_{B}T\right)}}{{1-\exp{\left(-\hbar\omega_{gf}/k_{B}T\right)}}}$
B $\frac{x_{2}-y_{2}}{x_{0}-x_{1}}$ $\frac{x_{1}-y_{1}}{y_{0}-x_{2}}$ $\frac{x_{0}-y_{0}}{y_{1}-y_{2}}$ $\frac{p_{e}-p_{f}}{p_{g}-p_{e}}$ $B=\frac{\exp{\left(-\hbar\omega_{ge}/k_{B}T\right)}-\exp{\left(-\hbar\omega_{gf}/k_{B}T\right)}}{{1-\exp{\left(-\hbar\omega_{ge}/k_{B}T\right)}}}$
C $\frac{x_{2}-y_{2}}{y_{0}-y_{1}}$ $\frac{x_{1}-y_{1}}{x_{0}-y_{2}}$ $\frac{x_{0}-y_{0}}{x_{1}-x_{2}}$ $\frac{p_{e}-p_{f}}{p_{g}-p_{f}}$ $C=\frac{\exp{\left(-\hbar\omega_{ge}/k_{B}T\right)}-\exp{\left(-\hbar\omega_{gf}/k_{B}T\right)}}{{1-\exp{\left(-\hbar\omega_{gf}/k_{B}T\right)}}}$
Appendix B Numerical Model
To describe the dynamics of the qubit population distribution, firstly we neglect effects of the readout resonator, and secondly, restrict ourselves by considering only the three lowest energy states of the transmon, |$\ket{g},\ \ket{e},\ \ket{f}$. We assume that the initial qubit state is thermal, which means that its density matrix is diagonal. Then we track the dynamics of only these diagonal terms, characterizing the population distribution, which significantly simplifies the calculations. We include only sequential decay and excitation processes with rates $\Gamma^{ge}_{\downarrow,\uparrow}$ and $\Gamma^{ef}_{\downarrow,\uparrow}$ for $g-e$ and $e-f$ transitions, respectively. Then the free time evolution of the qubit populations can be described by a system of differential equations:
$\begin{cases}\dot{p}_{f}&=p_{e}\Gamma^{ef}_{\uparrow}-p_{f}\Gamma^{ef}_{\downarrow},\\ \dot{p}_{e}&=p_{g}\Gamma^{ge}_{\uparrow}+p_{f}\Gamma^{ef}_{\downarrow}-p_{e}(\Gamma^{ge}_{\downarrow}+\Gamma^{ef}_{\uparrow}),\\ \dot{p}_{g}&=p_{e}\Gamma^{ge}_{\downarrow}-p_{g}\Gamma^{ge}_{\uparrow},\end{cases}$ (27)
where dotted functions denote time derivatives. The system has an analytical solution:
$p_{f}(t)$ $=\zeta_{f}e^{\alpha_{0}t}+\eta_{f}e^{\alpha_{1}t}+\xi_{f},$ (28)
$p_{e}(t)$ $=\zeta_{e}e^{\alpha_{0}t}+\eta_{e}e^{\alpha_{1}t}+\xi_{e},$ (29)
$p_{g}(t)$ $=\zeta_{g}e^{\alpha_{0}t}+\eta_{g}e^{\alpha_{1}t}+\xi_{g}.$ (30)
We find the following coefficients:
$\alpha_{0,1}=\frac{1}{2}$$\left\{-\left(\Gamma^{ef}_{\uparrow}+\Gamma^{ef}_{\downarrow}+\Gamma^{ge}_{\uparrow}+\Gamma^{ge}_{\downarrow}\right)\right.$ (31)
±$\pm\left[\left(\Gamma^{ef}_{\uparrow}+\Gamma^{ef}_{\downarrow}+\Gamma^{ge}_{\uparrow}+\Gamma^{ge}_{\downarrow}\right)^{2}\right.$
$-\left.\left.4\left(\Gamma^{ge}_{\downarrow}\Gamma^{ef}_{\downarrow}+\Gamma^{ge}_{\uparrow}\Gamma^{ef}_{\uparrow}+\Gamma^{ge}_{\uparrow}\Gamma^{ef}_{\downarrow}\right)\right]^{\frac{1}{2}}\right\},$
$\zeta_{f}$ $=\frac{\Gamma^{ef}_{\uparrow}\left(p_{e}^{(0)}-\xi_{e}\right)-\left(\Gamma^{ef}_{\downarrow}+\alpha_{1}\right)\left(p_{f}^{(0)}-\xi_{f}\right)}{\alpha_{0}-\alpha_{1}},$ (32)
$\zeta_{e}$ $=\frac{\alpha_{0}+\Gamma^{ef}_{\downarrow}}{\Gamma^{ef}_{\uparrow}}\zeta_{f},\hskip 14.22636pt\zeta_{g}=\frac{\Gamma^{ge}_{\downarrow}}{\Gamma^{ef}_{\uparrow}}\frac{\alpha_{0}+\Gamma^{ef}_{\downarrow}}{\alpha_{0}+\Gamma^{ge}_{\uparrow}}\zeta_{f},$ (33)
$\eta_{f}$ $=\frac{\left(\Gamma^{ef}_{\downarrow}+\alpha_{0}\right)\left(p_{f}^{(0)}-\xi_{f}\right)-\Gamma^{ef}_{\uparrow}\left(p_{e}^{(0)}-\xi_{e}\right)}{\alpha_{0}-\alpha_{1}},$ (34)
$\eta_{e}$ $=\frac{\alpha_{1}+\Gamma^{ef}_{\downarrow}}{\Gamma^{ef}_{\uparrow}}\eta_{f},\hskip 14.22636pt\eta_{g}=\frac{\Gamma^{ge}_{\downarrow}}{\Gamma^{ef}_{\uparrow}}\frac{\alpha_{1}+\Gamma^{ef}_{\downarrow}}{\alpha_{1}+\Gamma^{ge}_{\uparrow}}\eta_{f},$ (35)
[…]
where the canonical partition function $\mathcal{Z}=1+\Gamma^{ge}_{\uparrow}/\Gamma^{ge}_{\downarrow}+\Gamma^{ef}_{\uparrow}\Gamma^{ge}_{\uparrow}/\Gamma^{ef}_{\downarrow}\Gamma^{ge}_{\downarrow}$ imposes the normalization condition $p_{g}+p_{e}+p_{f}=1$.
The schematic representation of the model is presented in Fig. 1 c. We prepare the qubit with initial population $p_{g,e,f}^{(0)}$ and, for a fully thermalized qubit in the steady state ($t\rightarrow\infty$), observe residual populations $p_{i}^{(\infty)}=\xi_{i}$. To mimic state manipulation with microwave pulses, we define pulse matrices acting on a population vector $\vec{p}(t)=(p_{g}(t),p_{e}(t),p_{f}(t))^{T}$:
$M_{ge}$ $=\begin{pmatrix}1-\delta_{ge}&\delta_{ge}&0\\ \delta_{ge}&1-\delta_{ge}&0\\ 0&0&1\end{pmatrix},$
$M_{ef}$ $=\begin{pmatrix}1&0&0\\ 0&1-\delta_{ef}&\delta_{ef}\\ 0&\delta_{ef}&1-\delta_{ef}\end{pmatrix},$
where we introduce a phenomenological efficiency parameter $0\leq\delta\leq 1$ to simulate effects such as variation of pulse length and amplitude, detuning and finite fidelity. Unlike the real control pulses, these matrices act on the qubit population instantaneously. It is a realistic assumption if the pulse lengths are much smaller than the relaxation times. In this case the $\pi$-pulse duration corresponds to the time delay between application of corresponding instantaneous population transformations.
Figure 6: Comparison between the qubit population measurements and the numerical simulation for qubit R4-I. The simulation parameters are the same as in Fig. 3 b. The symbols show the experimental data and the lines correspond to the model.
We studied influence of the pulse efficiency on the error of the qubit temperature measurement (see Fig. 7). Our model predicts the relative error less than 10% at 200 mK, which is comparable with the experimental data.
Figure 7: Averaged relative error of the effective temperature depending on the $\pi_{ge,ef}$-pulse efficiency, based on the numerical model at the qubit thermalization temperature 200 mK. Apart from the variation in pulse efficiency, we use identical simulation parameters as in Fig. 3. The relative error is calculated by comparing temperatures extracted from imperfect pulses with finite efficiencies $\delta_{ge}$, $\delta_{ef}$, denoted as $T_{\delta_{ge},\delta_{ef}}$, and ideal pulses ($T_{\delta_{ge}=1,\delta_{ef}=1}$). We then average this quantity over all possible ways to extract qubit temperature $A_{1},A_{2},\dots,C_{3}$: $\langle[T_{\delta_{ge},\delta_{ef}}(\alpha)-T_{\delta_{ge}=1,\delta_{ef}=1}(\alpha)]/T_{\delta_{ge}=1,\delta_{ef}=1}(\alpha)\rangle_{\alpha=A_{1},A_{2},\dots,C_{3}}$.
Appendix C Error Estimation
C.1 Mutual Dependencies
As described in the main text, for thermometry we use temperature dependent functions $A,B$ and $C$ and each of them could be found in three different ways (see Table 2), which gives nine different values for temperature. Though the calculation of these parameters based on six independent $x_{j},\ y_{k}$, only six of these functions $A_{i},\ B_{i},\ C_{i}$ are mutually independent. That could be easily shown from relation $A_{i}\times B_{i}=C_{i}$, $i\in\{1,2,3\}$. That also leads to the mutual dependencies for the corresponding measurement errors. Indeed, one can get the same relation between the relative errors of $A,B$ and $C$:
$\left(\frac{\Delta C}{C}\right)^{2}=\left(\frac{\Delta A}{A}\right)^{2}+\left(\frac{\Delta B}{B}\right)^{2}.$ (37)
From this equation one can get the relation between the absolute errors:
$(\Delta C)^{2}=\left(\frac{C}{A}\right)^{2}(\Delta A)^{2}+\left(\frac{C}{B}\right)^{2}(\Delta B)^{2}.$ (38)
In the low temperature limit we assume $p_{f}=0$, $p_{e}\ll p_{g}\approx 1$. Using Eqs. (7), (8) we can get $C/B\approx 1$ and $C/A\approx p_{e}$, and as a result $(\Delta C)^{2}\approx(\Delta B)^{2}+p_{e}^{2}(\Delta A)^{2}$. If absolute errors for $A,B$ and $C$ are of the same order of magnitude (it is valid for most of the experimental results), absolute and relative errors for $B$ and $C$ are approximately equal: $(\Delta C)^{2}\approx(\Delta B)^{2}$ and $|\Delta C/C|\approx|\Delta B/B|$.
In order to find the relation between the errors of $A,B$ and $C$ and the measured temperature $T$ one can use logarithmic derivatives of the functions $A(T),B(T)$ and $C(T)$. In order to simplify the final expressions, in the low temperature limit expressions we can write
$\left|\frac{\Delta T}{T}\right|_{A}=\frac{e^{x}}{x}|\Delta A|=\frac{e^{x}-1}{x}\left|\frac{\Delta A}{A}\right|,$ (39)
$\left|\frac{\Delta T}{T}\right|_{B}=\frac{(e^{x}-1)^{2}}{xe^{x}}|\Delta B|=\frac{e^{x}-1}{xe^{x}}\left|\frac{\Delta B}{B}\right|,$ (40)
$\left|\frac{\Delta T}{T}\right|_{C}=\frac{e^{x}}{x}|\Delta C|=\frac{1}{x}\left|\frac{\Delta C}{C}\right|,$ (41)
where $x=\hbar\omega_{ge}/k_{B}T$. At low temperatures $x\rightarrow\infty$, Eqs. (40) and (41) give approximately equal errors, meaning that the relative errors of the temperatures $T_{B}$ and $T_{C}$ are similar in the relevant temperature range. That agrees very well with the data presented in this paper (see Sec. V and VI) and in Ref. [29].
Note, that although the coefficient between the relative error of $T$ and the relative error of $A$ in Eq. (39) is $e^{\hbar\omega_{ge}/k_{B}T}$ times higher than the same coefficients in Eqs. (40) and (41), the $A$ itself is $e^{\hbar\omega_{ge}/k_{B}T}$ times higher than $B$ and $C$, so it means that relative error for $T_{A}$ could be the same as for $T_{B}$ and $T_{C}$. The exact relation between errors for $T_{A}$ and for $T_{B}$ and $T_{C}$ depends on particular pure state responses and voltage noise, as will be shown in the next subsection.
C.2 Further Error Analysis
One can notice that columns 1, 2 and 3 in Table 2 have the same structure: $A,\ B$ and $C$ from one column could be written in a form $A=b/c$, $B=a/b$ and $C=a/c$, where $a,\ b$ and $c$ are the measured voltage differences. For example, in the first column $a=x_{2}-y_{2}$, $b=x_{0}-x_{1}$ and $c=y_{0}-y_{1}$. Also, one can write these voltage differences in the following form: $a=(p_{e}-p_{f})\Delta\varphi_{i}$, $b=(p_{g}-p_{e})\Delta\varphi_{i}$, $c=(p_{g}-p_{f})\Delta\varphi_{i}$. Here $\Delta\varphi_{i}=\varphi_{j}-\varphi_{k}$ for $i,j,k\in\{g,e,f\}$, $i\neq j\neq k$.
One can then make an additional assumption that the variance of a single measurement $x_{i},y_{i}$ is a sum of the voltage noise variance (defined by the measurement setup) and the variance of protective measurements with operator $\hat{M}$ of the qubit in a thermal state with density matrix $\hat{\rho}_{th}$: $\sigma_{x,y}^{2}=Tr(\hat{\rho}_{th}\hat{M}^{2})-Tr(\hat{\rho}_{th}\hat{M})^{2}$. Based on this, one can find absolute errors for measurement differences $a,b,c$:
$(\Delta a)^{2}$ = $=\ 2p_{e}p_{f}\Delta\varphi_{i}^{2}$
$+(p_{e}+p_{f})p_{g}(\Delta\varphi_{j}^{2}+\Delta\varphi_{k}^{2})+\sigma_{a}^{2},$ (42)
$(\Delta b)^{2}$ = $=\ 2p_{e}p_{g}\Delta\varphi_{i}^{2}$
$+(p_{e}+p_{g})p_{f}(\Delta\varphi_{j}^{2}+\Delta\varphi_{k}^{2})+\sigma_{b}^{2},$ (43)
$(\Delta c)^{2}$ $=\ (p_{g}+p_{e})p_{f}\Delta\varphi_{i}^{2}$
$+(p_{g}+p_{f})p_{e}(\Delta\varphi_{j}^{2}+\Delta\varphi_{k}^{2})+\sigma_{b}^{2},$ (44)
here $\sigma_{a}^{2},\sigma_{b}^{2},\sigma_{c}^{2}$ are the measurement voltage variances, which in general are not the same for different parameters $a,b,c$ and could be found from the measurements. Relative errors for the voltage differences $a,b$ and $c$ could be written as:
$\left(\frac{\Delta a}{a}\right)^{2}=\frac{2p_{e}p_{f}}{(p_{e}-p_{f})^{2}}$ $+\frac{(p_{e}+p_{f})p_{g}}{(p_{e}-p_{f})^{2}}F(\Delta\varphi_{i})$
$+$ $\frac{\sigma_{a}^{2}}{\Delta\varphi_{i}^{2}(p_{e}-p_{f})^{2}},$ (45)
$\left(\frac{\Delta b}{b}\right)^{2}=\frac{2p_{e}p_{g}}{(p_{g}-p_{e})^{2}}$ $+\frac{(p_{g}+p_{e})p_{f}}{(p_{g}-p_{e})^{2}}F(\Delta\varphi_{i})$
$+$ $\frac{\sigma_{b}^{2}}{\Delta\varphi_{i}^{2}(p_{g}-p_{e})^{2}},$ (46)
$\left(\frac{\Delta c}{c}\right)^{2}=\frac{(p_{e}+p_{g})p_{f}}{(p_{g}-p_{f})^{2}}$ $+\frac{(p_{g}+p_{f})p_{e}}{(p_{g}-p_{f})^{2}}F(\Delta\varphi_{i})$
$+$ $\frac{\sigma_{c}^{2}}{\Delta\varphi_{i}^{2}(p_{g}-p_{f})^{2}},$ (47)
where $F(\Delta\varphi_{i})=(\Delta\varphi_{j}^{2}+\Delta\varphi_{k}^{2})/\Delta\varphi_{i}^{2}$, $i\neq j\neq k$. Note, that function $F(\Delta\varphi_{i})$ reaches the minimum value of 0.5 when $\Delta\varphi_{i}=2\Delta\varphi_{j}=2\Delta\varphi_{k},\ i\neq j\neq k$. This gives a condition for the optimal difference of the pure state responses. Importantly, this condition can be satisfied only for the one out of three differences $\Delta\varphi_{i}$ at the same time.
In order to find relative errors for the $A,B,C$, one can use the following expressions:
$\left(\frac{\Delta A}{A}\right)^{2}=\left(\frac{\Delta b}{b}\right)^{2}+\left(\frac{\Delta c}{c}\right)^{2},$ (48)
$\left(\frac{\Delta B}{B}\right)^{2}=\left(\frac{\Delta a}{a}\right)^{2}+\left(\frac{\Delta b}{b}\right)^{2},$ (49)
$\left(\frac{\Delta C}{C}\right)^{2}=\left(\frac{\Delta a}{a}\right)^{2}+\left(\frac{\Delta c}{c}\right)^{2}.$ (50)
Using Eqs. (C.2, C.2, C.2) for relative errors for $a,b$ and $c$, one can substitute it to the equations above and get the final (but extremely bulky) result.
As in a previous subsection, one can find absolute errors for $A,B,C$ in the low-temperature limit ($p_{f}=0$, $p_{e}\ll p_{g}\approx 1$):
$(\Delta A)^{2}\approx 2p_{e}+F(\Delta\varphi_{i})+\frac{\sigma_{b}^{2}}{\Delta\varphi_{i}^{2}}+\frac{\sigma_{c}^{2}}{\Delta\varphi_{i}^{2}},$ (51)
$(\Delta B)^{2}\approx p_{e}F(\Delta\varphi_{i})+\frac{\sigma_{a}^{2}}{\Delta\varphi_{i}^{2}},$ (52)
$(\Delta C)^{2}\approx p_{e}F(\Delta\varphi_{i})+\frac{\sigma_{a}^{2}}{\Delta\varphi_{i}^{2}}.$ (53)
These equations illustrate several important features of errors of the functions $A,B$ and $C$. Firstly, the relative errors $\left|\Delta A/\langle A\rangle\right|,\left|\Delta B/\langle B\rangle\right|$ and $\left|\Delta C/\langle C\rangle\right|$ grow as $1/\sqrt{p_{e}}\propto\exp\left(\hbar\omega_{ge}/2k_{B}T\right)$ when $T\rightarrow 0$. Secondly, precision of the method could be limited by voltage noise, as it is for the measurements presented in this paper. Thirdly, in general, measurement precision highly depends on the differences $\Delta\varphi_{i}$ of the pure state responses. On one hand, it gives some room for the measurement protocol optimization in terms of maximization of the $\Delta\varphi_{i}^{2}$ and minimization of $F(\Delta\varphi_{i})$. But, unfortunately, this optimization should be done for each qubit separately. Finally, one can obtain the minimum possible values for the absolute errors in the low temperature limit: $(\Delta A)^{2}\approx 2p_{e}+0.5$, $(\Delta B)^{2}\approx 0.5p_{e}$, $(\Delta C)^{2}\approx 0.5p_{e}$.
To sum up, the precision of errors for the temperature estimations from $A,B$ and $C$ measurements are limited by voltage noise and strongly depend on the differences $\Delta\varphi_{i}$ of the pure state responses. So, in general, out of nine possible ways to extract temperature, there is no unique or universal answer as to which one is the optimal one. Instead, for every particular sample and measurement setup one needs to compare them and choose the best one.
Appendix D Sample Fabrication
[…]
Figure 8: Experimental setup. a) Room-temperature part of the setup, used in the experiment. b) Cryostat setup used for the low-temperature measurements. c) Decryption of the symbols, used in the panels a) and b), denoting instruments and other devices used in the measurement setup.
The next step was forming of Josephson junctions (JJs) (see Fig. 1 c). For that the sample was spincoated by a MMA/PMMA e-beam resist bilayer. A 800 nm thick layer of MMA EL-11 was applied by two steps of spincoating at 4000 rpm and baking at 160C for 2 min. Then a 200 nm layer of PMMA A4 was applied by spinning at 4000 rpm and baking at 160C for 10 min. Finally, an EBL step was performed to create a suspended mask for further JJ shadow deposition. The main and undercut doses for exposure of submicron structures was 950 $\mu$C/cm2 and 150 $\mu$C/cm2, respectively, with 1 nA beam current and 4 nm stepsize. The development of the mask was done in a two-step process: (i) development in MIBK:IPA 1:3 mixture for 25 s and (ii) in methylglycol:methanol 2:1 mixture for 15 s with a following rinsing in IPA for 30 s. After that the samples were loaded into an e-beam evaporator Plassys MEB700S2-III UHV for JJ deposition. First, the sample was pumped down to $5\cdot 10^{-7}$ mbar in the loadlock and an Ar milling was done to remove the resist residues at Ar pressure of $4\cdot 10^{-4}$ mbar with ion-beam gun (IBG) parameters of 60 mA beam current and 200 V beam voltage at angles ±$\pm 12^{\circ}$ for 30 s at each angle. Then the substrate was moved into the oxidation chamber and pumped to $1\cdot 10^{-7}$ mbar. After that the chambers were connected and Ti gettering was performed at a rate of 0.5 Å/s for 5 min. Then the first layer of 25 nm thick Al was deposited at angle $-12^{\circ}$ at 1 Å/s rate at pressure below $5\cdot 10^{-8}$ mbar in the oxidation chamber, after which the evaporation chamber was closed and an oxidation was done at a pressure of 18 mbar during 7 min. The second layer of 45 nm thick Al was deposited at $+12^{\circ}$ at 1 Å/s rate at a pressure below $1\cdot 10^{-7}$ mbar in the oxidation chamber. Finally, another oxidation was performed at pressure 20 mbar during 10 min. The lift-off procedure was done in hot acetone at $52^{\circ}$ for 3 h, then the sample was rinsed with IPA and dried with a N2 gun.
Now the third step of EBL was done to create patches (see Fig. 1 c) for a galvanic connection between JJs and the Nb ground plane. The MMA/PMMA e-beam resist bilayer was applied as described above and after that the sample was loaded into the EBL tool. The patch exposure was done with the dose of 950 $\mu$C/cm2 with 30 nA beam current and 20 nm stepsize. The resist was developed by putting into MIBK:IPA 1:3 mixture for 30 s and then into methylglycol:methanol 2:1 for 30 s followed by rinsing in IPA for 30 s. For the patch deposition we again used the Plassys tool. Analogously to the previous step, the sample in the loadlock was pumped to $5\cdot 10^{-7}$ mbar and a stronger Ar milling at zero angle was done for 2 min to remove the surface Nb and Al oxides. The Ar pressure was $4\cdot 10^{-4}$ mbar and the IBG parameters were the following: 120 mA beam current, 400 V beam voltage. Then the sample was moved to the oxidation chamber and pumped to $1\cdot 10^{-7}$ mbar, after which a deposition of 150 nm of Al was performed at the rate of 2 Å/s and pressure below $1\cdot 10^{-7}$ mbar in the oxidation chamber with consecutive oxidation at 20 mbar during 10 min. Another lift-off step was done similarly to the one described above.
For the low-temperature measurements the wafers were scribed with a diamond tip and cleaved into separate chips. The chips were mounted on a low-temperature sample holder using BF-6 glue and then wire-bonded using FS Bondtec 5330 with Al wire.